# Short-term plasticity recalls forgotten memories through a trampoline mechanism

### v1

@@ III Retrieval phase with short-term synaptic plasticity
When Eq. (14) is violated, then the terminal state of the network cannot be described by a unique fixed point within the basin of attraction of the condensed pattern. Multiple fixed-point solutions are possible, and their description requires more complex ansatz for the statistics of the noise $z$ [mezard1987, fischer1993spin].
Eqs. 9, 10, 11, 12, 13 and 14 describe the statistical properties of our recurrent neural network with short-term and long-term synaptic plasticity at a fixed point of the dynamics, and in particular, they allow to compute the average overlap with a target pattern. For $k=0$, we recover the equations obtained by Kühn for a Hopfield-RNN using the replica method [Kuhn1991statistical, Kuhn1993]. On the other hand, when $k\neq 0$, we see from Eq. (11) that turning short-term synaptic plasticity on leads to a shift of the screening constant $\Gamma$. As long as fixed points of the dynamics are considered, this shift can be equivalently obtained by setting $k=0$ and increasing the diagonal elements of the matrix $J_{ij}$ in Eq. (4) by a constant amount $kq$. Thus, within the validity region of the cavity method, short-term synaptic plasticity has the same effect as an additional self-coupling in a non-plastic network, with the salient aspect that the strength of the self-coupling needs to be self-consistently determined from the mean squared output $q$, given by Eq. (12).
The fixed point equations, Eqs. 9, 10, 11, 12, 13 and 14, can be solved numerically by iteration for different values of the fraction of patterns $\alpha$, output gain $\gamma$, and short-term synaptic strength $k$, by adapting the method of [Kuhn1993]. Once convergence is achieved, the value of the different order parameters can be measured, and a phase diagram for the properties of the fixed points of the dynamics of the network can be constructed, as a function of the fraction of stored patterns $\alpha$ and the inverse gain $\gamma^{-1}$, for different values of the short-term synaptic strength. These phase diagrams are illustrated in Fig. 2, which are obtained respectively without short-term synaptic plasticity, Fig. 2 (a), and for a moderate value of the plasticity strength, Fig. 2 (b).
@@
In both cases, four distinct phases can be identified. For low value of the load $\alpha$, and large enough values of the gain $\gamma$, a retrieval, single-fixed point phase (light green region in Fig. 2), the overlap of the network with the condensed pattern is close to unity, $m^{*}\approx 1$, and the stability condition given by Eq. (14) is obeyed. The attractors pertaining to each pattern are distinct from each other, and each of them is endowed by a unique fixed point. They are indeed global minima of the energy function $\mathcal{L}$ in Eq. (7). The dark green region in Fig. 2 represents instead a retrieval phase with multiple fixed points. Here too, we have $m^{*}\approx 1$, but the stability condition in Eq. (14) is violated. The basins of attractions around the patterns of the network are metastable, and they host multiple fixed points. Even though in this region the cavity calculation breaks down, previous works on associative memories [amit1985, amit1987, tokita1994replica] indicate that the cavity solution still yields an accurate description of the retrieval capabilities of the network. Upon increasing the fraction of patterns stored in the connectivity matrix, the system transitions toward a spin glass phase (orange region in Fig. 2), where the overlap with a condensed pattern is zero, $m^{*}=0$, and the stability condition given by Eq. (14) breaks down. Finally for low value of the fraction of stored patterns and high gains, the neural network has a unique fixed point, a paramagnetic state with zero overlap with the condensed pattern, $m^{*}=0$, and where the network is completely quiescent (blue region in Fig. 2).
@@
The salient result of our analysis is an understanding of the impact of short-term synaptic plasticity on the retrieval phase boundary, i.e. the curve $\alpha_{c}(\gamma,k)$ that separates the retrieval phases from the spin-glass phase. A comparison between Fig. 2 (a) and Fig. 2 (b) shows that short-term synaptic plasticity increases the retrieval region of the system at moderate gains $\gamma$, but yields little improvement at high gains. In particular, the maximal critical capacity that can be obtained from an Hopfield-RNN, $\alpha_{c}^{\text{max}}=\lim_{\gamma\to\infty}\alpha_{c}(\gamma,k=0)\approx 0.138$, is left unaffected by the presence of short-term synaptic plasticity. Indeed, the additional self-coupling term appearing in Eq. (11) becomes irrelevant in the limit $\gamma\to\infty$, as the output gain of the neuron converges to a sign function. On the other hand, the limit of infinite gain constitutes a strict upper bound to the critical capacity capacity. This follows from the boundedness of the neuron’s output function in the continuous-time framework [hopfield1984, Kuhn1993]: for any finite gain, $|\tanh(\gamma x)|<1$, whereas the infinite-gain limit yields a binary output $|\text{sgn}(x)|=1$. Consequently, the macroscopic overlap $m^{*}$ for finite $\gamma$ is strictly smaller than the one obtained in the infinite-gain limit. Such a reduction in the signal strength renders the retrieval state less robust against the interference noise from uncondensed patterns. Thus, we expect to have $\alpha_{c}(\gamma,k)<\alpha_{c}^{\text{max}}$ for finite $\gamma$.
@@
Inspection of Fig 2 (a-b) reveals that the the single-fixed point, retrieval region shrinks as short-term synaptic plasticity is turned on. Such a shrinkage can be understood by inspection of the stability condition in Eq. (14). A sufficient condition for the breakdown of the stability of the cavity method is obtained when the fixed point equation for the neuron output, Eq. (10) admits multiple solutions for some realization of the noise $z$, which results in a discontinuity in the derivative $\frac{\text{d}\phi}{\text{d}z}$. Such a situation corresponds to the opening of a gap in the distribution of the membrane potential of the population of neurons at the fixed point [Kuhn1993]. In particular, when $\sigma z=-\xi^{*}m^{*}$, Eq. (10) becomes $\phi=g(\Gamma\phi)$. If the addition of a self coupling yields higher values of the screening constant $\Gamma$, then it is easier to have a set of order parameters for which multiple solutions are possible. Moreover, since $g(x)=\tanh(\gamma x)$, we see how increasing the gain $\gamma$ makes it easier for the system to develop instabilities. Short-term synaptic plasticity thus increases the reactivity of the output of the population of neurons under small changes in their membrane potential, rendering the system more prone to fluctuations.
@@
We conclude this Section by commenting on the boundaries between the paramagnetic phase and the other phases in Fig. 2 (a-b). An analysis carried out in App. B.2 allows us to determine the critical gains $\gamma_{c}(k)$ and $\gamma_{g}(\alpha,k)$, denoting the boundaries between the paramagnetic phase and the retrieval phase as $\alpha\to 0$, and the boundary between the paramagnetic phase and the spin-glass phase at finite $\alpha$, respectively. Our result reads
@@
These two quantities are independent from $k$, and are, in particular, identical to their values in the absence of short-term synaptic plasticity. This situation is different from the case where a constant self-coupling term is added to the network, as studied in [Kuhn1993]. Such a feature is a consequence of the fact that, when short-term synaptic plasticity is turned on, the strength of the self-coupling needs to be self-consistently determined from the network activity, according to Eq. (12). In the paramagnetic state, the network is quiescent, $\phi=0$, and therefore also the self-overlap $q$ vanishes, $q=0$, which explain the absence of a dependence from $k$ in Eq. (15).
@@ IV Dynamical plastic retrieval
To understand the retrieval ability of the network when the static cavity method breaks down, we study the dynamics of the network when its capacity exceeds the critical threshold, by employing a combination of finite size simulations and dynamical mean field theory (DMFT) [sompolinsky1988, clark2024dynamic, clark2025transient].
@@ IV.1 Dynamical mean-field theory
The dynamical mean field theory framework allows to derive an effective equation for the dynamics of a tagged neuron in the network, by integrating out the influence of all the other neurons and the plastic synapses onto the tagged neuron. The Markovian, many-body dynamics in Eqs. 1, 2, 3, 4 and 5 becomes an effective non-Markovian dynamics for an individual neuron. The effect of the surrounding system is encoded into a memory term and a stochastic noise. In the limit $N\to\infty$, thanks to the fully connected topology of the model, the statistics of the noise becomes Gaussian and colored. Moreover, both the noise autocorrelation function and the memory kernel can be self-consistently expressed in terms of the statistics of the effective single-neuron process. The dynamical evolution of network population averages, such as the overlap with a target pattern, can be obtained as averages over the effective single-neuron process under different realizations of the effective noise.
The detailed derivation of the DMFT is presented in App. C. The final result is an effective single-site stochastic process describing the evolution of the membrane potential $x(t)$ of a tagged neuron in the system, which reads
@@
The first term on the right hand side stems from the single-site decay term each neuron is subjected to. The term $\xi^{*}m^{*}(t)$ involves the dynamics overlap $m(t)$ of the network with a condensed pattern, and it plays the role of an effective signal driving the neuron toward the target pattern. The dynamics overlap is determined self-consistently as
@@
Such a kernel propagates in time the activity of a neuron in the network from a past time $t^{\prime}$ to a later time $t$, provided that the neuron is typically firing at both these times. The strength of the signal provided by the plastic kernel decays exponentially over the short-term plasticity time-scale $p$.
Eqs. 16, 17, 18, 19, 20, 21 and 22 describe the DMFT of a neuron in our recurrent neural network with short-term synaptic plasticity. Their validity extends across the whole $\gamma^{-1}-\alpha$ phase diagram, and can thus be used to describe the dynamical evolution of the network, even in the catastrophic forgetting phase. For $k=0$, we recover the case with purely long-term associative memory [clark2025transient]. In the presence of short-term synaptic plasticity, the plastic kernel $K_{\text{plast}}$ strengthens the memory term in Eq. (16), propagating the activity of the neuron from the past. This is the dynamical realization of the trampoline-like effect that was commented at the beginning of Sec. III, when discussing the energy landscape of the recurrent neural network. The structure of Eqs. 16, 17, 18, 19, 20, 21 and 22 is quite complex, and even in the absence of short-term synaptic plasticity, an analytical solution is not known. We can, however, make progress by solving Eqs. 16, 17, 18, 19, 20, 21 and 22 numerically. The procedure is detailed in App. D. In particular, the numerical solution allows us to track the dynamical evolution of the overlap $m^{*}(t)$, when the network is initialized in the vicinity of a condensed pattern, $m^{*}(0)\neq 0$. Such a nonzero initial overlap value is achieved by tuning the distribution of the initial condition $x(0)$. Such a numerical solution is accompanied by finite-size simulations of the dynamics of a full neural-network (details in App. E), which we use to support the validity of the numerical integration of the DMFT equations.
@@
The capacity is chosen above the maximal critical capacity of the system $\alpha>\alpha_{c}^{\text{max}}\approx 0.138$, so that, from the perspective of the static cavity method, the network is in a spin-glass, forgetting phase. For $k=0$, the overlap $m^{*}(t)$ initially increases, reaching a maximum of height $m^{*}(t_{\text{max}})=0.82$ around a time $t_{\text{max}}\approx 1.5$, while at longer times, the overlap $m^{*}(t)$ slowly decays down to 0. The nonmonotonic behavior of $m^{*}(t)$ for $k=0$ is known as transient retrieval, and it is an indication of the complex structure of the high-dimensional basins of attraction of the Hopfield recurrent neural network above the critical capacity. The slow decay at longer times is instead a well-recognized instance of catastrophic forgetting, and it demonstrates the inability of the network to robustly store patterns above its critical capacity (see Fig 6 in App. E for results from finite-size simulations at longer time in the absence of synaptic plasticity). However, in the presence of short-term synaptic plasticity, the situation becomes quite different. For small values of $k$, we still observe the phenomena of transient retrieval and catastrophic forgetting, albeit the maximum value of the overlap achieved during transient retrieval increases with $k$, and the decay of the overlap occurs more slowly. However, when $k$ exceeds a critical value, which for the parameter chosen in this case is $\approx 0.8$, the overlap $m^{*}(t)$ plateaus to a nonzero value at large times. Interestingly, the value of the long-time plateau varies nonmonotonically with the strength of the short-term synaptic plasticity, as shown in Fig. 3 (b). Moreover, the value of the long-time plateau of $m^{*}(t)$ for large values of $k$ can even be smaller than the maximum overlap value achieved in the absence of short-term synaptic plasticity. These findings demonstrate that short-term synaptic plasticity can stabilize the dynamic transient retrieval phase of a Hopfield recurrent neural network, generating new metastable states which considerably overlap with the stored patterns. These metastable states then trap the network in the vicinity of the target pattern, allowing for robust retrieval. Short-term synaptic plasticity allows past network activity to be propagated forward in time to the current neural activity, providing a memory effects that slow down the network and facilitates its arrest, hindering the escape from the target pattern and the access to the catastrophic forgetting regime. In the trampoline-like analogy put forward at the beginning of Sec. III, the continuous reshaping by short-term synaptic plasticity of the effective neural energy landscape leads to a slowdown of the network, in the same way a ball rolling through a trampoline is hindered in its motion with respect to a ball rolling over a solid surface.

### v2

@@ III Retrieval phase with short-term synaptic plasticity
When Eq. (14) is violated, then the terminal state of the network cannot be described by a unique fixed point within the basin of attraction of the condensed pattern. Multiple fixed-point solutions are possible, and their description requires more complex ansatz for the statistics of the noise $z$ [39, 14].
Eqs. 9, 10, 11, 12, 13 and 14 describe the statistical properties of our recurrent neural network with short-term and long-term synaptic plasticity at a fixed point of the dynamics, and in particular, they allow to compute the average overlap with a target pattern. For $k=0$, we recover the equations obtained by Kühn for a Hopfield-RNN using the replica method [31, 30]. On the other hand, when $k\neq 0$, we see from Eq. (11) that turning short-term synaptic plasticity on leads to a shift of the screening constant $\Gamma$. As long as fixed points of the dynamics are considered, this shift can equivalently be obtained by setting $k=0$ and increasing the diagonal elements of the matrix $J_{ij}$ in Eq. (4) by a constant amount $kq$. Thus, within the validity region of the cavity method, short-term synaptic plasticity has the same effect as an additional self-coupling in a non-plastic network, with the salient aspect that the strength of the self-coupling needs to be self-consistently determined from the mean squared output $q$, given by Eq. (12).
The fixed-point equations, Eqs. 9, 10, 11, 12, 13 and 14, can be solved numerically by iteration for different values of the fraction of patterns $\alpha$, output gain $\gamma$, and short-term synaptic strength $k$, by adapting the method of [30]. Once convergence is achieved, the value of the different order parameters can be measured, and a phase diagram for the properties of the fixed points of the dynamics of the network can be constructed, as a function of the fraction of stored patterns $\alpha$ and the inverse gain $\gamma^{-1}$, for different values of the short-term synaptic strength. These phase diagrams are illustrated in Fig. 2, which are obtained respectively without short-term synaptic plasticity, Fig. 2 (a), and for a moderate value of the plasticity strength, Fig. 2 (b).
@@
In both cases, four distinct phases can be identified. For low value of the load $\alpha$, and large enough values of the gain $\gamma$, a retrieval, single-fixed point phase (light green region in Fig. 2), the overlap of the network with the condensed pattern is close to unity, $m^{*}\approx 1$, and the stability condition given by Eq. (14) is obeyed. The attractors pertaining to each pattern are distinct from each other, and each of them is endowed by a unique fixed point. They are indeed global minima of the energy function $\mathcal{L}$ in Eq. (7). The dark green region in Fig. 2 instead represents a retrieval phase with multiple fixed points. Here too, we have $m^{*}\approx 1$, but the stability condition in Eq. (14) is violated. The basins of attractions around the patterns of the network are metastable, and they host multiple fixed points. Even though in this region the cavity calculation breaks down, previous works on associative memories [3, 4, 51] indicate that the cavity solution still yields an accurate description of the retrieval capabilities of the network. Upon increasing the fraction of patterns stored in the connectivity matrix, the system transitions toward a spin glass phase (orange region in Fig. 2), where the overlap with a condensed pattern is zero, $m^{*}=0$, and the stability condition given by Eq. (14) breaks down. Finally for low value of the fraction of stored patterns and high gains, the neural network has a unique fixed point, a paramagnetic state with zero overlap with the condensed pattern, $m^{*}=0$, and where the network is completely quiescent (blue region in Fig. 2).
@@
The salient result of our analysis is an understanding of the impact of short-term synaptic plasticity on the retrieval phase boundary, i.e. the curve $\alpha_{c}(\gamma,k)$ that separates the retrieval phases from the spin-glass phase. A comparison between Fig. 2 (a) and Fig. 2 (b) shows that short-term synaptic plasticity increases the retrieval region of the system at moderate gains $\gamma$, but yields little improvement at high gains. In particular, the maximal critical capacity that can be obtained from an Hopfield-RNN, $\alpha_{c}^{\text{max}}=\lim_{\gamma\to\infty}\alpha_{c}(\gamma,k=0)\approx 0.138$, is left unaffected by the presence of short-term synaptic plasticity. Indeed, the additional self-coupling term appearing in Eq. (11) becomes irrelevant in the limit $\gamma\to\infty$, as the output gain of the neuron converges to a sign function. On the other hand, the limit of infinite gain constitutes a strict upper bound to the critical capacity capacity. This follows from the boundedness of the neuron’s output function in the continuous-time framework [24, 30]: for any finite gain, $|\tanh(\gamma x)|<1$, whereas the infinite-gain limit yields a binary output $|\text{sgn}(x)|=1$. Consequently, the macroscopic overlap $m^{*}$ for finite $\gamma$ is strictly smaller than the one obtained in the infinite-gain limit. Such a reduction in the signal strength renders the retrieval state less robust against the interference noise from uncondensed patterns. Thus, we expect to have $\alpha_{c}(\gamma,k)<\alpha_{c}^{\text{max}}$ for finite $\gamma$.
@@
Inspection of Fig 2 (a-b) reveals that the the single-fixed point, retrieval region shrinks as short-term synaptic plasticity is turned on. Such shrinkage can be understood by inspection of the stability condition in Eq. (14). A sufficient condition for the breakdown of the stability of the cavity method is obtained when the fixed point equation for the neuron output, Eq. (10) admits multiple solutions for some realization of the noise $z$, which results in a discontinuity in the derivative $\frac{\text{d}\phi}{\text{d}z}$. Such a situation corresponds to the opening of a gap in the distribution of the membrane potential of the population of neurons at the fixed point [30]. In particular, when $\sigma z=-\xi^{*}m^{*}$, Eq. (10) becomes $\phi=g(\Gamma\phi)$. If the addition of a self coupling yields higher values of the screening constant $\Gamma$, then it is easier to have a set of order parameters for which multiple solutions are possible. Moreover, since $g(x)=\tanh(\gamma x)$, we see how increasing the gain $\gamma$ makes it easier for the system to develop instabilities. Short-term synaptic plasticity thus increases the reactivity of the output of the population of neurons under small changes in their membrane potential, rendering the system more prone to fluctuations.
@@
We conclude this Section by commenting on the boundaries between the paramagnetic phase and the other phases in Fig. 2 (a–b). An analysis carried out in App. C.2 allows us to determine the critical gains $\gamma_{c}(k)$ and $\gamma_{g}(\alpha,k)$, denoting the boundaries between the paramagnetic phase and the retrieval phase as $\alpha\to 0$, and the boundary between the paramagnetic phase and the spin-glass phase at finite $\alpha$, respectively. Our result reads
@@
These two quantities are independent of $k$, and are, in particular, identical to their values in the absence of short-term synaptic plasticity. This situation is different from the case where a constant self-coupling term is added to the network, as studied in [30]. This feature is a consequence of the fact that, when short-term synaptic plasticity is turned on, the strength of the self-coupling needs to be self-consistently determined from the network activity, according to Eq. (12). In the paramagnetic state, the network is quiescent, $\phi=0$, and therefore the self-overlap $q$ also vanishes, $q=0$, which explains the absence of dependence on $k$ in Eq. (15).
@@ IV Dynamical plastic retrieval
To understand the retrieval ability of the network when the static cavity method breaks down, we study the dynamics of the network when its capacity exceeds the critical threshold, employing a combination of finite-size simulations and dynamical mean-field theory (DMFT) [49, 50, 48, 7, 8].
@@ IV.1 Dynamical mean-field theory
The dynamical mean field theory framework allows us to derive an effective equation for the dynamics of a tagged neuron in the network, by integrating out the influence of all the other neurons and the plastic synapses onto the tagged neuron. The Markovian, many-body dynamics in Eqs. 1, 2, 3, 4 and 5 becomes an effective non-Markovian dynamics for an individual neuron. The effect of the surrounding system is encoded in a memory term and a stochastic noise. In the limit $N\to\infty$, thanks to the fully connected topology of the model, the statistics of the noise become Gaussian and colored. Moreover, both the noise autocorrelation function and the memory kernel can be self-consistently expressed in terms of the statistics of the effective single-neuron process. The dynamical evolution of network population averages, such as the overlap with a target pattern, can be obtained as averages over the effective single-neuron process under different realizations of the effective noise.
The detailed derivation of the DMFT is presented in App. D. The final result is an effective single-site stochastic process describing the evolution of the membrane potential $x(t)$ of a tagged neuron in the system, which reads
@@
The first term on the right-hand side stems from the single-site decay term each neuron is subjected to. The term $\xi^{*}m^{*}(t)$ involves the dynamical overlap $m(t)$ of the network with a condensed pattern, and it plays the role of an effective signal driving the neuron toward the target pattern. The dynamics overlap is determined self-consistently as
@@
Such a kernel propagates the activity of a neuron in the network forward in time, from a past time $t^{\prime}$ to a later time $t$, provided that the neuron is typically firing at both these times. The strength of the signal provided by the plastic kernel decays exponentially over the short-term plasticity time-scale $p$.
Eqs. 16, 17, 18, 19, 20, 21 and 22 describe the DMFT of a neuron in our recurrent neural network with short-term synaptic plasticity. Their validity extends across the whole $\gamma^{-1}-\alpha$ phase diagram, and can thus be used to describe the dynamical evolution of the network, even in the catastrophic forgetting phase. For $k=0$, we recover the case with purely long-term associative memory [8]. In the presence of short-term synaptic plasticity, the plastic kernel $K_{\text{plast}}$ strengthens the memory term in Eq. (16), propagating the activity of the neuron from the past. This is the dynamical realization of the trampoline-like effect that was commented at the beginning of Sec. III, when discussing the energy landscape of the recurrent neural network. The structure of Eqs. 16, 17, 18, 19, 20, 21 and 22 is quite complex, and even in the absence of short-term synaptic plasticity, an analytical solution is not known. We can, however, make progress by solving Eqs. 16, 17, 18, 19, 20, 21 and 22 numerically. The procedure is detailed in App. E. In particular, the numerical solution allows us to track the dynamical evolution of the overlap $m^{*}(t)$, when the network is initialized in the vicinity of a condensed pattern, $m^{*}(0)\neq 0$. Such a nonzero initial overlap value is achieved by tuning the distribution of the initial condition $x(0)$. Such a numerical solution is accompanied by finite-size simulations of the dynamics of a full neural-network (details in App. F), which we use to support the validity of the numerical integration of the DMFT equations.
@@
The capacity is chosen above the maximal critical capacity of the system $\alpha>\alpha_{c}^{\text{max}}\approx 0.138$, so that, from the perspective of the static cavity method, the network is in a spin-glass, forgetting phase. For $k=0$, the overlap $m^{*}(t)$ initially increases, reaching a maximum of height $m^{*}(t_{\text{max}})=0.82$ around a time $t_{\text{max}}\approx 1.5$, while at longer times, the overlap $m^{*}(t)$ slowly decays down to 0. The nonmonotonic behavior of $m^{*}(t)$ for $k=0$ is known as transient retrieval, and it is an indication of the complex structure of the high-dimensional basins of attraction of the Hopfield recurrent neural network above the critical capacity. The slow decay at longer times is instead a well-recognized instance of catastrophic forgetting, and it demonstrates the inability of the network to robustly store patterns above its critical capacity (see Fig 6 (a) in App. F for results from finite-size simulations at longer times). However, in the presence of short-term synaptic plasticity, the situation becomes quite different. For small values of $k$, we still observe the phenomena of transient retrieval and catastrophic forgetting, albeit the maximum value of the overlap achieved during transient retrieval increases with $k$, and the decay of the overlap occurs more slowly. However, when $k$ exceeds a critical value, which for the parameter chosen in this case is $\approx 0.8$, the overlap $m^{*}(t)$ plateaus to a nonzero value at large times. Interestingly, the value of the long-time plateau varies nonmonotonically with the strength of the short-term synaptic plasticity, as shown in Fig. 3 (b). Moreover, the value of the long-time plateau of $m^{*}(t)$ for large values of $k$ can even be smaller than the maximum overlap value achieved in the absence of short-term synaptic plasticity (see Fig 6 in App. F for results from finite-size simulations at longer times, both with and without short-term synaptic plasticity). These findings demonstrate that short-term synaptic plasticity can stabilize the dynamic transient retrieval phase of a Hopfield recurrent neural network, generating new metastable states which considerably overlap with the stored patterns. These metastable states then trap the network in the vicinity of the target pattern, allowing for robust retrieval. Short-term synaptic plasticity allows past network activity to be propagated forward in time to the current neural activity, providing a memory effects that slow down the network and facilitates its arrest, hindering the escape from the target pattern and the access to the catastrophic forgetting regime. In the trampoline-like analogy put forward at the beginning of Sec. III, the continuous reshaping by short-term synaptic plasticity of the effective neural energy landscape leads to a slowdown of the network, in the same way a ball rolling through a trampoline is hindered in its motion with respect to a ball rolling over a solid surface.